Trail Conditions

January 21, 2022 -24º

Yesterday Kyle groomed 52 North to the headpond and then Dave hopped in and groomed 515, 520, 521, 522, 523, and 52 South to the CI Road, it was 3:00 am when he finished. We will be out on the trails again this morning. Allain removed the busted hoses from the 2001 BR180 and had them made at Sunny Corner Enterprises, got them back on late yesterday, and with any luck today we will have the system filled and bled, so should be back on the trail by late today. The 2009 Br180 is waiting for parts then the track will be fixed and reinstalled, looks like this will be next week. 515 on the Neguac side is now groomed but the markers haven’t been put out yet. Like and share draw will be Monday, so get sharing. Get Out and Enjoy YOUR trails.